Two grams of musk oil are required for each bottle of Mink Caress, a very popular perfume made by a small company in western Siberia. The cost of the musk oil is $2.10 per gram. Budgeted production of Mink Caress is given below by quarters for Year 2 and for the first quarter of Year 3:

    

 

Year 2

 

Year 3

 
  First Second Third Fourth   First  
  Budgeted production, in bottles 84,000 114,000 174,000 124,000   94,000  
 

    

    Musk oil has become so popular as a perfume ingredient that it has become necessary to carry large inventories as a precaution against stock-outs. For this reason, the inventory of musk oil at the end of a quarter must be equal to 20% of the following quarter’s production needs. Some 33,600 grams of musk oil will be on hand to start the first quarter of Year 2.

  

Required:

Prepare a direct materials budget for musk oil, by quarter and in total, for Year 2. (Round "Unit cost of raw materials" answers to 2 decimal places.)

 
 
Mink Caress
Direct Materials Budget - Year 2
  Quarter  
  First Second Third Fourth Year
Required production in units of finished goods          
Units of raw materials needed per unit of finished goods          
Units of raw materials needed to meet production          
           
Total units of raw materials needed

Units of raw materials to be purchased

Unit cost of raw materials

Cost of raw materials to purchased
Answer #1
Concepts and reason

Material Purchase budget: This budget can be interpreted as the budget which shows the opening inventory, quantity of direct material purchased during the year in units and with its associated cost. This budget is one of the components of the master budget. Closing inventory is deducted in the computation of the purchased units in the material purchase units.

Budget: A budget is a financial plan made by the company for a definite period such as for months or years. This is generally expressed in the monetary terms. Budget is generally prepared for making the crucial decisions about the future.

Fundamentals

Inventory: It is goods of the company which can be in forms of raw material, work in progress and finished goods. The company held these goods for sale in the ordinary course of business it shows the capability of the company to manufactured goods in what quantity it is showed in the asset side of the balance sheet.

Direct materials: These are the materials which are used in the Production process which constitutes an integral portion of the final product, the cost of which is identifiable and traceable directly to it.

 

Equivalent Units of Production: It is the term used in the system of cost accounting so as to arrive at the per unit cost of any product. This term is linked with that number of units which are not yet completed at the end of the particular period.

Closing inventory: it refers to the amount of goods stored that a business has left at the end of the accounting period. The account of closing stock is credited because these are the goods that remain unsold and the amount is to be calculated at the end of the fiscal or accounting year.

Opening inventory: It is the value of the inventory which is there with the firm at the start of the accounting period; it is also the ending inventory of the previous period.

Compute total units of raw material needed of the company for a year, using MS-excel as shown below:

Working notes:

Compute units of raw material needed to meet production for the first quarter, using the equation as shown below:

Unitsofrawmaterial=(Requiredproductioninunits×Perunitofrawmaterialneeded)=84,000×2=168,000\begin{array}{c}\\{\rm{Units of raw material}} = \left( \begin{array}{l}\\{\rm{Required production in units}}\\\\ \times {\rm{Per unit of raw material needed}}\\\end{array} \right)\\\\ = 84,000 \times 2\\\\ = 168,000\\\end{array}

Hence, the units of raw material needed to meet production are 168,000

Compute units of raw material needed to meet production for the second quarter, using the equation as shown below:

Unitsofrawmaterial=(Requiredproductioninunits×Perunitofrawmaterialneeded)=114,000×2=228,000\begin{array}{c}\\{\rm{Units of raw material}} = \left( \begin{array}{l}\\{\rm{Required production in units}}\\\\ \times {\rm{Per unit of raw material needed}}\\\end{array} \right)\\\\ = 114,000 \times 2\\\\ = 228,000\\\end{array}

Hence, the units of raw material needed to meet production are 228,000

Compute units of raw material needed to meet production for the third quarter, using the equation as shown below:

Unitsofrawmaterial=(Requiredproductioninunits×Perunitofrawmaterialneeded)=174,000×2=348,000\begin{array}{c}\\{\rm{Units of raw material}} = \left( \begin{array}{l}\\{\rm{Required production in units}}\\\\ \times {\rm{Per unit of raw material needed}}\\\end{array} \right)\\\\ = 174,000 \times 2\\\\ = 348,000\\\end{array}

Hence, the units of raw material needed to meet production are 348,000

Compute units of raw material needed to meet production for the fourth quarter, using the equation as shown below:

Unitsofrawmaterial=(Requiredproductioninunits×Perunitofrawmaterialneeded)=124,000×2=248,000\begin{array}{c}\\{\rm{Units of raw material}} = \left( \begin{array}{l}\\{\rm{Required production in units}}\\\\ \times {\rm{Per unit of raw material needed}}\\\end{array} \right)\\\\ = 124,000 \times 2\\\\ = 248,000\\\end{array}

Hence, the units of raw material needed to meet production are 248,000

Compute units of raw material needed to meet production for the year, using the equation as shown below:

Unitsofrawmaterial=(Requiredproductioninunits×Perunitofrawmaterialneeded)=496,000×2=992,000\begin{array}{c}\\{\rm{Units of raw material}} = \left( \begin{array}{l}\\{\rm{Required production in units}}\\\\ \times {\rm{Per unit of raw material needed}}\\\end{array} \right)\\\\ = 496,000 \times 2\\\\ = 992,000\\\end{array}

Hence, the units of raw material needed to meet production are 992,000

Compute required production in units of finished goods for a year, using the equation shown below:

Requiredproductioninunits=(Requiredproductionunitsoffirstquarter+secondquarter+thirdquarter+fourthquarter)=84,000+114,000+174,000+124,000=496,000\begin{array}{c}\\{\rm{Required production in units}} = \left( \begin{array}{l}\\{\rm{Required production units of first quarter}}\\\\ + {\rm{second quarter}}\\\\ + {\rm{third quarter}} + {\rm{fourth quarter}}\\\end{array} \right)\\\\ = 84,000 + 114,000 + 174,000 + 124,000\\\\ = 496,000\\\end{array}

Hence, the required production in units for the year is 496,000 units

Compute ending raw material of first quarter using the equation below:

Endingrawmaterialinventory=Unitsofrawmaterialofsecondquarter×%=228,000×20%=45,600\begin{array}{c}\\{\rm{Ending raw material inventory}} = {\rm{Units of raw material of second quarter}} \times \% \\\\ = 228,000 \times 20\% \\\\ = 45,600\\\end{array}

Hence, the ending raw material inventory is 45,600

Compute ending raw material of second quarter using the equation below:

Endingrawmaterialinventory=Unitsofrawmaterialofthirdquarter×%=348,000×20%=69,600\begin{array}{c}\\{\rm{Ending raw material inventory}} = {\rm{Units of raw material of third quarter}} \times \% \\\\ = 348,000 \times 20\% \\\\ = 69,600\\\end{array}

Hence, the ending raw material inventory is 69,600

Compute ending raw material of third quarter using the equation below:

Endingrawmaterialinventory=Unitsofrawmaterialoffourthquarter×%=248,000×20%=49,600\begin{array}{c}\\{\rm{Ending raw material inventory}} = {\rm{Units of raw material of fourth quarter}} \times \% \\\\ = 248,000 \times 20\% \\\\ = 49,600\\\end{array}

Hence, the ending raw material inventory is 49,600

Compute ending raw material of fourth quarter using the equation below:

Endingrawmaterialinventory=Unitsofrawmaterialofyear3×2×%=94,000×2×20%=37,600\begin{array}{c}\\{\rm{Ending raw material inventory}} = {\rm{Units of raw material of year 3}} \times {\rm{2}} \times \% \\\\ = 94,000 \times 2 \times 20\% \\\\ = 37,600\\\end{array}

Hence, the ending raw material inventory is 37,600

Prepare a direct material budget of the company for the year, using MS-Excel as shown below:

Compute units of raw material to be purchased for the first quarter by using the equation below:

Unitsofrawmaterialtobepurchased=(TotalunitsofrawmaterialBeginninginventory)=213,60033,600=180,000\begin{array}{c}\\{\rm{Units of raw material to be purchased}} = \left( \begin{array}{l}\\{\rm{Total units of raw material}}\\\\ - {\rm{Beginning inventory}}\\\end{array} \right)\\\\ = 213,600 - 33,600\\\\ = 180,000\\\end{array}

Hence, the units of raw material to be purchased is 180,000

Compute units of raw material to be purchased for the second quarter by using the equation below:

Unitsofrawmaterialtobepurchased=(TotalunitsofrawmaterialBeginninginventory)=297,60045,600=252,000\begin{array}{c}\\{\rm{Units of raw material to be purchased}} = \left( \begin{array}{l}\\{\rm{Total units of raw material}}\\\\ - {\rm{Beginning inventory}}\\\end{array} \right)\\\\ = 297,600 - 45,600\\\\ = 252,000\\\end{array}

Hence, the units of raw material to be purchased is 252,000

Compute units of raw material to be purchased for the third quarter by using the equation below:

Unitsofrawmaterialtobepurchased=(TotalunitsofrawmaterialBeginninginventory)=397,60069,600=328,000\begin{array}{c}\\{\rm{Units of raw material to be purchased}} = \left( \begin{array}{l}\\{\rm{Total units of raw material}}\\\\ - {\rm{Beginning inventory}}\\\end{array} \right)\\\\ = 397,600 - 69,600\\\\ = 328,000\\\end{array}

Hence, the units of raw material to be purchased is 328,000

Compute units of raw material to be purchased for the fourth quarter by using the equation below:

Unitsofrawmaterialtobepurchased=(TotalunitsofrawmaterialBeginninginventory)=285,60049,600=236,000\begin{array}{c}\\{\rm{Units of raw material to be purchased}} = \left( \begin{array}{l}\\{\rm{Total units of raw material}}\\\\ - {\rm{Beginning inventory}}\\\end{array} \right)\\\\ = 285,600 - 49,600\\\\ = 236,000\\\end{array}

Hence, the units of raw material to be purchased is 236,000

Compute units of raw material to be purchased for the year by using the equation below:

Unitsofrawmaterialtobepurchased=(TotalunitsofrawmaterialBeginninginventory)=1,029,60033,600=996,000\begin{array}{c}\\{\rm{Units of raw material to be purchased}} = \left( \begin{array}{l}\\{\rm{Total units of raw material}}\\\\ - {\rm{Beginning inventory}}\\\end{array} \right)\\\\ = 1,029,600 - 33,600\\\\ = 996,000\\\end{array}

Hence, the units of raw material to be purchased is 996,000

Compute the cost of raw material to be purchased for the first quarter by using the equation below:

Costofrawmaterialtobepurchased=(Unitsofrawmaterial×Unitcostofrawmaterial)=180,000×2.1=$378,000\begin{array}{c}\\{\rm{Cost of raw material to be purchased}} = \left( \begin{array}{l}\\{\rm{Units of raw material}}\\\\ \times {\rm{Unit cost of raw material}}\\\end{array} \right)\\\\ = 180,000 \times 2.1\\\\ = \$ 378,000\\\end{array}

Hence, the cost of raw material to be purchased is $378,000

Compute the cost of raw material to be purchased for the second quarter by using the equation below:

Costofrawmaterialtobepurchased=(Unitsofrawmaterial×Unitcostofrawmaterial)=252,000×2.1=$529,200\begin{array}{c}\\{\rm{Cost of raw material to be purchased}} = \left( \begin{array}{l}\\{\rm{Units of raw material}}\\\\ \times {\rm{Unit cost of raw material}}\\\end{array} \right)\\\\ = 252,000 \times 2.1\\\\ = \$ 529,200\\\end{array}

Hence, the cost of raw material to be purchased is $529,200

Compute the cost of raw material to be purchased for the third quarter by using the equation below:

Costofrawmaterialtobepurchased=(Unitsofrawmaterial×Unitcostofrawmaterial)=328,000×2.1=$688,800\begin{array}{c}\\{\rm{Cost of raw material to be purchased}} = \left( \begin{array}{l}\\{\rm{Units of raw material}}\\\\ \times {\rm{Unit cost of raw material}}\\\end{array} \right)\\\\ = 328,000 \times 2.1\\\\ = \$ 688,800\\\end{array}

Hence, the cost of raw material to be purchased is $688,800.

Compute the cost of raw material to be purchased for four quarter by using the equation below:

Costofrawmaterialtobepurchased=(Unitsofrawmaterial×Unitcostofrawmaterial)=236,000×2.1=$495,600\begin{array}{c}\\{\rm{Cost of raw material to be purchased}} = \left( \begin{array}{l}\\{\rm{Units of raw material}}\\\\ \times {\rm{Unit cost of raw material}}\\\end{array} \right)\\\\ = 236,000 \times 2.1\\\\ = \$ 495,600\\\end{array}

Hence, the cost of raw material to be purchased is $495,600

Compute the cost of raw material to be purchased for the year by using the equation below:

Costofrawmaterialtobepurchased=(Unitsofrawmaterial×Unitcostofrawmaterial)=996,000×2.1=$2,091,600\begin{array}{c}\\{\rm{Cost of raw material to be purchased}} = \left( \begin{array}{l}\\{\rm{Units of raw material}}\\\\ \times {\rm{Unit cost of raw material}}\\\end{array} \right)\\\\ = 996,000 \times 2.1\\\\ = \$ 2,091,600\\\end{array}

Hence, the cost of raw material to be purchased is $2,091,600

Ans: